OpenStudy (anonymous):

Solve the equation algebraically: tanxsinxcosx - 1 = 0 Now, can someone explain why the answer has no solution? Haalllppppp

OpenStudy (anonymous):

perhaps if we rewrite \[\tan(x)\cos(x)\sin(x)\] it will be more clear

OpenStudy (anonymous):

since \[\tan(x)=\frac{\sin(x)}{\cos(x)}\] you are looking at \[\frac{\sin(x)}{\cos(x)}\sin(x)\cos(x)=\sin^2(x)\]

OpenStudy (anonymous):

solving \[\sin^2(x)-1=0\] is not hard what do you get?

OpenStudy (anonymous):

So we don't factor it like \[(\sin(x)-1) (\sin(x)+1) = 0?\]

OpenStudy (anonymous):

you can if you like

OpenStudy (anonymous):

or you can just write \[\sin^2(x)=1\]so \[\sin(x)=1\] or \[\sin(x)=-1\]

OpenStudy (anonymous):

idk apparently the answer doesn't have a solution, that's where i got confused :/

OpenStudy (anonymous):

we are not done yet!

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

once we solve \[\sin(x)=1\] for \(x\) we can see why there is no solution did you solve it yet?

OpenStudy (anonymous):

How so? I thought the solution can also be \[\pi/2\] or \[3\pi/2\]

OpenStudy (anonymous):

so that makes it undefined?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

since at those number, cosine is zero

OpenStudy (anonymous):

don't forget you stared with \[\tan(x)\sin(x)\cos(x)\] and since cosine is zero at those number, a) tangent is not defined there and b) cosine is zero, so you would actually have \(\frac{1}{0}\times 1\times 0\) moral of the story you cannot cancel zeros

OpenStudy (anonymous):

ahah, that makes a lot of sense now hehe :) tysm!